Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  WTSC_{Entrant #18}
Round 1 Due Date : [July 27]  WTSC_{Entrant #19}
Round 1 Due Date : [July 27]  WTSC_{Entrant #20}
Round 1 Due Date : [July 27]  WTSC_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  EARC_{Entrant #18}
Round 1 Due Date : [July 27]  EARC_{Entrant #19}
Round 1 Due Date : [July 27]  EARC_{Entrant #20}
Round 1 Due Date : [July 27]  EARC_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  Comber_{Entrant #18}
Round 1 Due Date : [July 27]  Comber_{Entrant #19}
Round 1 Due Date : [July 27]  Comber_{Entrant #20}
Round 1 Due Date : [July 27]  Comber_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  FSC_{Entrant #18}
Round 1 Due Date : [July 27]  FSC_{Entrant #19}
Round 1 Due Date : [July 27]  FSC_{Entrant #20}
Round 1 Due Date : [July 27]  FSC_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  RRPC_{Entrant #18}
Round 1 Due Date : [July 27]  RRPC_{Entrant #19}
Round 1 Due Date : [July 27]  RRPC_{Entrant #20}
Round 1 Due Date : [July 27]  RRPC_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC_{Entrant #18}
Round 1 Due Date : [July 27]  DURC_{Entrant #19}
Round 1 Due Date : [July 27]  DURC_{Entrant #20}
Round 1 Due Date : [July 27]  DURC_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #1}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #2}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #3}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #4}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #5}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #6}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #7}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #8}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #9}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #10}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #11}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #12}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #13}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #14}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #15}

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #16}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #17}

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC_{Entrant #18}
Round 1 Due Date : [July 27]  DRC_{Entrant #19}
Round 1 Due Date : [July 27]  DRC_{Entrant #20}
Round 1 Due Date : [July 27]  DRC_{Entrant #21}

Place behind first bull of target

Place behind first bull of target

Place behind first bull of target Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target

Trinity Top Ten 2001

Place behind first bull of target